OpenStudy (anonymous):

the difference between a number and 5 when multiplied by 3 and 36 what is the number

OpenStudy (anonymous):

(x-5)*3 = 36?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

3x-15=36 3x=36+15 x=51/3 x=17